Accounting Rate of Return

WeCare Clinic is planning on investing in some new echocardiogram equipment that will require an initial outlay of $150,000. The system has an expected life of five years and no expected salvage value. The investment is expected to produce the following net cash flows over its life: $82,000, $85,000, $80,000, $81,000, and $94,000.

Required:

1. Calculate the annual net income for each of the five years.

Net Income
Year 1 $
Year 2 $
Year 3 $
Year 4 $
Year 5 $

2. Calculate the accounting rate of return. Enter your answer as a whole percentage value (for example, 16% should be entered as "16").
%

3. What if a second competing revenue-producing investment has the same initial outlay and salvage value but the following cash flows (in chronological sequence): $94,000, $94,000, $94,000, $82,000, and $29,000? Calculate its accounting rate of return. Enter your answer as a whole percentage value (for example, 16% should be entered as "16").
%

Using the accounting rate of return metric, which project should be selected: the first or the second?

Why might the second project be preferred over the first project?
Answer #1
1
Net Income
Year 1 82000-30000 = 52000
Year 2 85000-30000 = 55000
Year 3 80000-30000 = 50000
Year 4 81000-30000 = 51000
Year 5 94000-30000 = 64000
Average Net Income = 52000 + 55000 + 50000 + 51000 + 64000
5
= 272000
5
= 54400
Average Investment = 150000 + 0
2
= 75000
2
Accounting Rate of return = Average Net Income
Average Investment
= 54400
75000
= 72.50%
= 72.5

3)

Net Income
Year 1 94000-30000 = 64000
Year 2 94000-30000 = 64000
Year 3 94000-30000 = 64000
Year 4 82000-30000 = 52000
Year 5 29000-30000 = -1000
Average Net Income = 64000 64000 64000 52000 -1000
5
= 243000
5
= 48600
Average Investment = 150000 + 0
2
= 75000
Accounting Rate of return = Average Net Income
Average Investment
= 48600
75000
= 64.80%
= 64.8

Basis accounting rate of return first project should be opted.

However, the second project may be opted considering the more inflows in initial years.

if we consider time value of money and pay back period, then option 2 is more likely to be adopted